OpenStudy (anonymous):

Simplify. -7xy^2/28x^5y

OpenStudy (anonymous):

\[\frac{ x^a }{ x^b } = x^{a-b}\]

OpenStudy (anonymous):

\[\frac{ -7 }{ 28 }x^{1-5}y^{2-1}\]

OpenStudy (anonymous):

\[= \frac{ -1 }{ 4 } x^{-4}y = -\frac{ y }{ 4x^4 }\]